please help me figure out this question.
The perimeter of a rectangular swimming pool is not to exceed 70 ft. The
length is to be twice the width. what widths, w, will meet these conditions?
thank you for helping me out(:

Answer by stanbon(75887)

The perimeter of a rectangular swimming pool is not to exceed 70 ft. The
length is to be twice the width. what widths, w, will meet these conditions?
-----
P <= 70 ft.
L = 2W
-----
Inequality:
2(L+W) <= 70
L+W <= 35
------
2W+W <= 35
3W <= 35
Width <= 11 2/3 feet
----
Length = 2W <= 23 1/3 feet
==============================
Cheers,
Stan H.
==============